# Feeding the Crave: How People with Eating Disorders Get Trapped in the Perpetual Cycle of Digital Food Content

### v3

@@ A.1. Exploratory Study Participants Demographics
Table 4. Exploratory study participants demographic, eating disorder information, and stated digital food content. The listed items (i.e., food content and food delivery apps) in the stated digital food content are identified from open-ended responses in surveys and interviews. BED refers to Binge Eating Disorder and BN stands for Bulimia Nervosa. BED is characterized by recurrent episodes of excessive eating in a short timeframe (e.g., 2 hours) and feeling a loss of control over their eating behavior (Dingemans et al., 2002). BN involves a cycle of binge eating and compensatory behaviors, such as self-induced vomiting (Fairburn and Garner, 1986).
A.2. In-Depth Study Participants Demographics
@@
Table 5. In-depth study participants demographic, ED information, and stated digital food content. The stated digital food content items are identified from open-ended interview responses. BED refers to Binge Eating Disorder and BN stands for Bulimia Nervosa. BED is characterized by recurrent episodes of excessive eating in a short timeframe (e.g., 2 hours) and feeling a loss of control over their eating behavior (Dingemans et al., 2002). BN involves a cycle of binge eating and compensatory behaviors, such as self-induced vomiting (Fairburn and Garner, 1986).
A.3. In-Depth Study Semi-structured Interview Protocol - Digital Food Content Consumption and ED Symptoms
This is an English-translated summary of the key questions we aimed to address. In the semi-structured interviews, we delved deeper and investigated topics beyond these specific questions.
Warm Up
Hello, and thank you once again for participating in this user study. As mentioned in the email, today’s interview is scheduled to last up to an hour. You are welcome to turn your camera on or off as you prefer. If there are any questions during the interview that you do not want to answer, please feel free to decline. With your consent, we will now begin recording the interview.
Exposures to digital food content and eating disorder symptoms
First, I will ask you some questions about the relationship between digital food content and the eating disorder symptoms you are experiencing.
@@ A.3. In-Depth Study Semi-structured Interview Protocol - Digital Food Content Consumption and ED Symptoms
    Do you think that your digital food content consumption is associated with your ED symptoms? Why do you think it is (not) related?
@@
    What kinds of food content do you think are related?
    For each content,
    1. (a)
      What motivates you to consume the content?
    2. (b)
      Could you share any specific instances/experiences where your eating disorder symptoms led to the consumption of the content?
    3. (c)
      Are there any other factors, besides eating disorder symptoms, that drive you to consume the content?
    4. (d)
      Could you share any specific instances/experiences where consuming the content impacted your eating disorder symptoms?
    5. (e)
      Have you tried regulating your consumption of the content? If so, why and how?
Digital supports to regulate digital food content consumption

### v4

@@ 3.1. Participants
Table 1. Exploratory study participants’ demographic, eating disorder information, and stated digital food content. The listed items (i.e., food content and food delivery apps) in the stated digital food content are identified from open-ended responses in surveys and interviews. BED refers to Binge Eating Disorder and BN stands for Bulimia Nervosa. BED is characterized by recurrent episodes of excessive eating in a short timeframe (e.g., 2 hours) and feeling a loss of control over their eating behavior (Dingemans et al., 2002). BN involves a cycle of binge eating and compensatory behaviors, such as self-induced vomiting (Fairburn and Garner, 1986).
We recruited 23 participants (aged 18-43, mean=26.2 years; 19 females, 1 male, 1 non-binary, 2 chose not to disclose) through an advertisement post on an online community for people with EDs under the permission of the moderators in South Korea (cop., 2018). The participant gender composition is predominantly female, reflecting universal ED demographics (87.4% of identified South Korean binge eating disorder patients being female (Service, 2023), and threefold higher lifetime prevalence of binge eating disorder and bulimia nervosa in females in the US (Hudson et al., 2007; Galmiche et al., 2019)). Table 1 provides an overview of participant demographics, ED information, and types of digital food content mentioned by each participant. Our study intentionally included individuals who were either formally diagnosed or self-diagnosed, as has been done in prior research (Choi et al., 2024; Feuston et al., 2020; Choi et al., 2025). By doing so, we aimed to encompass a population that may not have sought professional treatment, providing insights into the often overlooked experiences of these individuals within a clinical context.
3.2. Study Protocol
The study involved an online survey to investigate how people with EDs associate digital food content with their ED symptoms. Additionally, we conducted remote semi-structured interviews to gain a deeper understanding. We employed the Eating Disorder Examination Questionnaire 6.0 (EDE-Q) to assess if our participants represented the eating disorder population (Fairburn and Beglin, 2008). [^1] Higher scores on the EDE-Q indicate problematic eating behaviors and attitudes. The EDE-Q scores of our study’s interviewees (4.56±0.56) were generally higher than those of individuals diagnosed with ED (4.02±1.28) (Aardoom et al., 2012). To address the social stigma associated with EDs (Puhl and Suh, 2015), participants were given the option to opt out of any part of the study, which consists of a survey, an interview, and the EDE-Q, they find uncomfortable.
In surveys, we intentionally refrained from introducing digital food content to accurately assess how participants perceived its influence on their symptoms. Instead, we asked about their computing device usage associated with disordered eating compared with normal meals. Specifically, survey questions included are: Are there any differences in your digital device usage before/during/after eating disorder episodes than without eating disorder episodes? If so, which device(s) do you use, and for what purpose? This approach allowed us to identify unique device usages, including consumption of digital food content, associated with their disordered eating.
Each Zoom interview, lasting approximately thirty minutes to one hour, was recorded and transcribed after the interview with consent. Given the stigma surrounding EDs (Puhl and Suh, 2015), we allowed the participants to freely turn off their cameras during the interview. In the interviews, participants shared their experiences regarding how various digital device usages are associated with and result in disordered eating. We also inquired about any apps and content they perceived to affect their ED symptoms. For participants who did not respond to the survey, we also asked the survey questions during the interview. We did not bring up digital food content unless participants voluntarily mentioned it, which is consistent with our survey approach. Only the interviewees, not survey-only respondents, were rewarded approximately USD 23 for a one-hour interview. The study materials, including survey questions, interview questions and participants’ responses, were administered in Korean and subsequently translated into English. All translations were reviewed by the authors to ensure accuracy and preserve the integrity of participants’ original statements.
3.3. Analysis
3.4. Results
Sixteen out of twenty-three participants voluntarily linked digital food content to their disordered eating behaviors, revealing a nuanced relationship. Our analysis identified two primary types of content associated with such behaviors: food media and food delivery apps. Specifically, eleven participants mentioned food media, including recipe blogs, food photos, chewing or crunching ASMR videos, [^2] An Autonomous Sensory Meridian Response (ASMR) refers to a tingling sensation in response to specific visual, auditory, or tactile stimuli. ASMR videos often feature soothing voices, tapping, eating, and other soothing sounds. and eating broadcasts. Nine participants referred to food delivery apps. These findings reaffirm the connection between digital food content and unhealthy dietary habits (Bodenlos and Wormuth, 2013; Spence et al., 2019).
Participants’ narratives revealed that food media played an ambivalent role in their experiences with disordered eating, serving both as a self-regulation strategy and a trigger for binge episodes. Several participants described watching food-related videos in an effort to suppress the urge to binge eat or to gain emotional comfort. For example, P15 shared, “I watched Mukbangs to try to stop myself from binge eating to get some vicarious satisfaction. I watch them to suppress the urge.” She also noted that she “sometimes watches videos of others who binge and regret it, especially when feeling stressed. It gives me a sense of solidarity and reminds me that I’m not alone, and that recovery is still possible by seeing people in such videos who make up their minds again even after such relapses.” Likewise, P3 shared, “I’ve tried watching slow eating videos while eating, to help myself pace and prevent binge episodes.
Yet, the same form of media often triggered harmful outcomes as emphasized in the previous study (Choi et al., 2024). Some participants, including P3 and P15, who described food media as a helpful coping tool, also reported that it sometimes led to binge eating. P7 explained, “When I repeatedly view such things (suggested food media), I unconsciously feel the urge to eat it. It would be okay if I could eat it without feeling guilty about it, but if I, even once, feel guilty (because it would lead to weight gain), it piles up on and on, pushing me to binge eat. So, in the long run, it does have an impact.
Through our exploratory study, we identified the ambivalent impact of food media—both supporting and undermining recovery efforts within a single engagement. To further understand the underlying dynamics of this ambivalence, we conducted a follow-up in-depth interview study focusing on (1) individuals’ motivations for engaging with digital food content, (2) the conflicting impacts such content has on their eating disorder experiences, and (3) the strategies they adopt to avoid exposure to triggering content. This deeper investigation aims to inform the design of digital interventions that acknowledge both the harms and the personally meaningful benefits embedded in everyday interactions with food content among individuals with eating disorders.
4. In-depth Study
4.1. Participants
@@
Table 2. In-depth study participants’ demographics, ED information, and stated digital food content. The stated digital food content items are identified from open-ended interview responses.
We recruited 22 participants (aged 18-41, mean=26.7 years; 21 females and 1 male) through advertisement posts on the online social support communities for people with EDs under the permission of the moderators in South Korea (cop., 2018; KakaoTalk, 2022). Table 2 presents participants’ demographics and ED diagnosis information. Two in-depth study participants, P9 and P11 (P6 and P8 in the exploratory study, respectively), also participated in our exploratory interviews. To be eligible for the study, participants must (1) be over 18 years old and (2) have an ED. Since many people do not seek treatment for their EDs, eligibility for this study, as in our exploratory study, was not contingent upon a diagnosis. Participants, however, had to identify themselves as having ED symptoms (e.g., restriction of food intake, binge eating, or compensatory behaviors). Again, by recruiting users who did not seek treatment or had discontinued it, we could shed light on the needs of users whose struggles are largely ignored in a clinical setting.
4.2. Study Protocol
All phases of our study were IRB-approved and conducted remotely due to the high social stigma of EDs (Puhl and Suh, 2015). Participants responded to a survey via email before the interview. The survey included the EDE-Q and questions about their demographics and ED symptoms. [^3] The average EDE-Q score of the participants was 4.40±1.09, compared with averages of individuals diagnosed with ED (4.02±1.28) (Aardoom et al., 2012). Following this, we conducted semi-structured interviews with all participants on Zoom, each lasting approximately thirty minutes to one hour. We allowed participants to turn off their cameras. Every interview was recorded and transcribed with consent from the participants. Our interview delved into the motivations and practices of digital food content consumption by asking questions about participants’ experiences. We encouraged participants to share their varied experiences with different types of food media, covering various food videos, pictures, and posts. The detailed interview protocol is in our Supplementary Materials. The study materials, including interview questions and participants’ responses, were administered in Korean and subsequently translated into English. All translations were reviewed by the authors to ensure accuracy and preserve the integrity of participants’ original statements.
4.3. Analysis
We conducted the inductive thematic analysis (Braun and Clarke, 2012) of the interview responses to investigate why and how individuals with EDs interact with digital food content and how such content relates to disordered eating behaviors. After transcribing the interviews, the first and second authors individually looked through all transcriptions to fully understand the interview content. They independently generated initial codes and, through iterative discussions on emerging themes, addressed inconsistencies, resolved disagreements, and ensured that the data strongly supported the identified themes. To finalize the analysis, all authors participated in comprehensive discussions. Once the analysis was completed, each theme was titled and described.
5. In-Depth Study Results
Building on our exploratory study’s observation of impact ambivalence—the coexistence of recovery-supporting and ED-exacerbating effects within the same engagement—we structured our in-depth analysis to examine how this ambivalence manifests across participants’ motivations, perceived impacts, and avoidance practices.
5.1. Patterns of Digital Food Content Consumption
Figure 1. A flowchart describing interaction patterns with digital food content. The OR symbol ($\oplus$) denotes that the process continues in multiple branches. The Summoning Junction symbol ($\otimes$) indicates that multiple branches converge into a single process.
[…]
Figure 1 is the flowchart that depicts patterns of digital food content engagement. The detailed explanation of these patterns is described in Section 5.1.
Before delving into specific motivations, impacts, and avoidance practices, we outline the overarching patterns of participants’ engagement with digital food content. This overview provides context for how these different aspects unfold over time and interact with one another (Figure 1).
Across the interviews, participants described their engagement as a recurring sequence:
@@ 5.1. Patterns of Digital Food Content Consumption
    Engage with Digital Food Content: Participants began consuming digital food content, often by chance or with specific motivations, such as seeking comfort, managing urges, or browsing out of habit.
@@
    Recovery‑Supporting Impact: Early in their engagement, they felt the content satisfied their motivations and did not recognize harmful effects.